solve for x: (3x-2)(x-1)=0
Answer by stanbon(75887)

solve for x: (3x-2)(x-1)=0
---
3x-2 = 0 or x-1 = 0
x = 2/3 or x = 1
=============================
Cheers,
Stan H.